### v1

# Short geodesics and multiplicities of eigenvalues of hyperbolic surfaces

Abstract.
In this paper, we obtain upper bounds on the multiplicity of Laplacian eigenvalues for closed hyperbolic surfaces in terms of the number of short closed geodesics and the genus $g$. For example, we show that if the number of short closed geodesics is sublinear in $g$, then the multiplicity of the first eigenvalue is also sublinear in $g$. This makes new progress on a conjecture by Colin de Verdière in the mid 1980s.
1. Introduction
Let $X_{g}$ be a closed hyperbolic surface of genus $g\geq 2$, and let $\Delta$ denote the Laplacian on $X_{g}$. By classical results in spectral theory, the spectrum of $\Delta$ consists of a sequence of discrete nonnegative eigenvalues counted with multiplicities:
$0=\lambda_{0}(X_{g})<\lambda_{1}(X_{g})\leq\lambda_{2}(X_{g})\leq\cdots\to+\infty.$
For any $\lambda>0$, denote by $m(\lambda)$ the multiplicity of $\lambda$ in the spectrum of $\Delta$. In this paper, we are interested in studying the upper bounds of their multiplicities, especially in $m(\lambda_{1})$.
For any constant $\epsilon>0$ that may depend on $g$, define
$\mathcal{N}_{\epsilon}(X_{g}):=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\},$
and set $N_{\epsilon}(X_{g})=\#{\mathcal{N}_{\epsilon}(X_{g})}$. Our first result is as follows.
Theorem 1.1.
Let $X_{g}$ be a closed hyperbolic surface of genus $g$. Then there exists a uniform constant $K\geq 1$ such that for any constant $\epsilon\in(0,1)$, the multiplicity of $\lambda_{1}(X_{g})$ satisfies
$m(\lambda_{1}(X_{g}))\leq\frac{K}{\epsilon^{2}}\cdot\frac{g}{\log\log\left(\frac{10g}{N_{\epsilon}(X_{g})+1}\right)}.$
Remark.
  1. (1)
    If $N_{\epsilon_{0}}(X_{g})\leq f(g)=o(g)$ for some $\epsilon_{0}>0$ and positive function $f:\mathbb{Z}^{\geq 2}\to\mathbb{Z}$ with sublinear growth, then Theorem 1.1 tells that
    $m(\lambda_{1}(X_{g}))\ll\frac{g}{\log\log\left(\frac{g}{f(g)}\right)}.$
    This upper bound is new and also sublinear in $g$. In particular, if $f(g)\ll g^{\alpha}$ for any fixed $\alpha\in(0,1)$, then we have
    $m(\lambda_{1}(X_{g}))\ll\frac{g}{\log\log g}.$
    See Figure 1 for an example.
  2. (2)
    If $N_{\epsilon_{0}}(X_{g})=0$ for some $\epsilon_{0}>0$, which is equivalent to saying that $X_{g}$ has injectivity radius $\geq\epsilon_{0}$, as above Theorem 1.1 tells that
    $m(\lambda_{1}(X_{g}))\ll\frac{g}{\log\log g}.$
    This was due to Letrouit-Machado [LM24] for this case. We remark here that Jiang-Tidor-Yao-Zhang-Zhao [JTY+21] proved the same sublinear upper bound for the second largest eigenvalue of connected $g$-vertex graphs with bounded maximum degree.
  3. (3)
    By the classical Collar Lemma, it is known that $N_{\epsilon}(X_{g})\leq 3g-3$ whenever $\epsilon$ is small. Thus, if we choose $\epsilon=0.5$, Theorem 1.1 then tells that for any closed hyperbolic surface $X_{g}$ of genus $g$,
    $m(\lambda_{1}(X_{g}))\ll g.$
    The linear upper bound was first obtained by Besson [Bes80].
Figure 1. A closed hyperbolic surface $\mathcal{X}_{g}$ of genus $g$, where $\gamma_{i}\in\mathcal{N}_{\epsilon_{0}}(\mathcal{X}_{g})$ and $N_{\epsilon_{0}}(\mathcal{X}_{g})=o(g)$ for some fixed $\epsilon_{0}>0$. So $m(\lambda_{1}(\mathcal{X}_{g}))=o(g)$.
A key novel ingredient in proving Theorem 1.1 is to find a linear subspace $S$ of the eigenspace $E(\lambda_{1}(X_{g}))$ with respect to $\lambda_{1}(X_{g})$ such that the dimension $\dim(S)$ of the subspace $S$ is uniformly comparable to $\dim(E(\lambda_{1}(X_{g})))=m_{1}(\lambda_{1}(X_{g}))$, and the $L^{2}$-norm of each eigenfunction in $S$ on $X_{g}$ is uniformly comparable to its $L^{2}$-norm on the $\epsilon$-thick part of $X_{g}$. And then restricted to the thick part of $X_{g}$, we follow a strategy similar to that in the recent works [JTY+21, LM24] to complete the proof of Theorem 1.1.
Our second result concerns the multiplicities of small eigenvalues. For any $\delta\in(0,1/2)$, set
$\epsilon(\delta)=\operatorname{arcsinh}\frac{1}{\sinh(1/\delta+2)},$
and for any $\epsilon\in(0,\epsilon(\delta))$, set
$I_{\epsilon}(X_{g})=\textit{the number of components of }X_{g}\setminus\mathcal{N}_{\epsilon}(X_{g}).$
Our second theorem is as follows.
Theorem 1.2.
Let $X_{g}$ be a closed hyperbolic surface of genus $g$, and let $\lambda>0$ be an eigenvalue of $X_{g}$. If $\lambda\leq 1/4-\delta^{2}$, then for any $\epsilon\in(0,\epsilon(\delta))$, there exists a constant $C(\epsilon)>0$ only depending on $\epsilon$ such that
$m(\lambda)\leq C(\epsilon)\sqrt{\lambda}\cdot g+24I_{\epsilon}(X_{g}).$
Remark.
  1. (1)
    If $\lambda_{1}(X_{g})=o(1)$ and $I_{\epsilon_{0}}(X_{g})=o(g)$ for some fixed $\epsilon_{0}$ as $g\to\infty$, then Theorem 1.2 tells that $m(\lambda_{1}(X_{g}))=o(g)$. Note that $N_{\epsilon_{0}}(X_{g})$ may have growth rate $g$, in which case Theorem 1.1 only gives a linear upper bound for $m(\lambda_{1}(X_{g}))$. See Figure 2 for an example.
  2. (2)
    Suppose $\epsilon\ll 1/(\log g)^{\kappa}$ for some $\kappa>0$ as $g\to\infty$. If $\mathcal{N}_{\epsilon}(X_{g})$ separates $X_{g}$, and $I_{\epsilon_{0}}(X_{g})\ll g/(\log g)^{\kappa}$ for some fixed $\epsilon_{0}$, then the upper bound given by Theorem 1.2 is slightly stronger than that given by Theorem 1.1.
Figure 2. A closed hyperbolic surface $\mathcal{Y}_{g}$ of genus $g$ separated by $g+1$ geodesics into two pieces, where $\gamma_{i}\in\mathcal{N}_{\epsilon}(\mathcal{Y}_{g})$ and $\epsilon=o(1)$ as $g\to\infty$. So $m(\lambda_{1}(\mathcal{Y}_{g}))=o(g)$.
We remark here that for a family of closed hyperbolic surfaces $\{X_{g}^{t}\}_{t\in(0,1]}$ of genus $g$ that has $(3g-3)$ pairwise disjoint simple closed geodesics simultaneously pinching to zero as $t\to 0$, it is known from e.g., [SWY80] or [HW25, Proof of Theorem 21] that $\lambda_{2g-3}(X_{g}^{t})\to 0$ uniformly as $t\to 0$. However, both Theorem 1.1 and Theorem 1.2 cannot distinguish the first $(2g-3)$ nonzero eigenvalues of this particular family $\{X_{g}^{t}\}_{t\in(0,1]}$. To our knowledge, it is not even known whether the multiplicity $m(\lambda_{1}(X_{g}^{t}))$ is sublinear in $g$ when $t$ is close to 00.
Related works
The first upper bound on the multiplicity of Laplacian eigenvalues on closed Riemannian surfaces was established by Cheng [Che76], who showed that for any closed Riemannian surface of genus $g$, the multiplicity of the $i$-th eigenvalue $\lambda_{i}$ satisfies
$m(\lambda_{i})\leq\frac{1}{2}(2g+i+1)(2g+i+2),\qquad\operatorname{\forall}i\geq 1.$
This was later improved by Besson [Bes80], who obtained the first linear upper bound:
$m(\lambda_{i})\leq 4g+2i+1.$
Many refinements of Besson’s bound have since been obtained. Nadirashvili [Nad88] improved the bound for $g\geq 2$ to $14g+2i-$ 1; Sévennec [Sév02] improved it further for $g\geq 2$, $i=1$ to $32g+$ 3; and Bourque–Petri [FBP23] improved the bound for $i=1$ to $12g-$ 1 for hyperbolic surfaces of sufficiently large genus.
Colin de Verdière’s conjecture
In [CdV86], Colin de Verdière conjectured that the multiplicity of the first nonzero Laplacian eigenvalue on a closed orientable surface $M_{g}$ of genus $g$ is bounded from above by
$\mathrm{chr}(M_{g})-1,$
where chromatic number $\mathrm{chr}(M_{g})$ is the supremum of the natural numbers $n$ for which the complete graph on $n$ vertices can be embedded in $M_{g}$. It is known from [RY68] of Ringel-Youngs that
$\mathrm{chr}(M_{g})=\Big{\lfloor}\frac{1}{2}\big{(}7+\sqrt{48g+1}\big{)}\Big{\rfloor}.$
In support of this, Colbois and Colin de Verdière constructed, in [CCdV88], for every $g\geq 3$, a hyperbolic surface $X_{g}$ such that
$m(\lambda_{1}(X_{g}))=\Big{\lfloor}\frac{1+\sqrt{8g+1}}{2}\Big{\rfloor},$
demonstrating that the maximal multiplicity grows at least on the order of $\sqrt{g}$. However, counterexamples to the conjecture are found for $g=10$ and $g=17$ by Bourque-Gruda Mediavilla-Petri-Pineault in [BGMPP23].
Nevertheless, for closed hyperbolic surfaces of high genus, it is quite interesting to determine the right growth rate of the maximal multiplicity of the first nonzero eigenvalue of the Laplacian. Is it $\sqrt{g}$ in the above conjecture of Colin de Verdière, or $\frac{g}{\log\log g}$ as shown above, or something else? Several sublinear upper bounds on the multiplicity of the first nonzero eigenvalue of the Laplacian on hyperbolic surfaces have been established under certain assumptions [LM24, FBP23].
Notations.
[…]
Plan of the paper
In Section 2 we introduce the standard thick-thin decomposition of closed hyperbolic surfaces. In Section 3 we use the number of short closed geodesics to construct a subspace of $\lambda$-eigenspace on which each eigenvalue has $L^{2}$-norm concentrated in the thick part, which will be applied in the proofs of Theorem 1.1 and Theorem 1.2. In Section 4 we prove Theorem 1.1. And we prove Theorem 1.2 in Section 5.
Acknowledgements
We would like to thank all the participants in our seminar at Tsinghua University on Teichmüller theory for helpful discussions on this project. The first named author is supported by the China Postdoctoral Science Foundation No. 2024M761591. The second named author is partially supported by NSFC grants No. 12171263, 12361141813, and 12425107.
2. Thick-thin decomposition
Let $X_{g}$ be a closed hyperbolic surface of genus $g$, and let $\operatorname{dist}(\cdot,\cdot)$ denote the distance function on $X_{g}$. For any simple closed geodesic $\gamma\subset X_{g}$, the collar around $\gamma$ of width $w$ is defined by
$\operatorname{\mathcal{C}}(\gamma,w):=\{x\in X_{g}|\ \operatorname{dist}(x,\gamma)\leq w\}.$
Let $\ell(\gamma)$ be the length of $\gamma$. The Collar Lemma asserts that the collar of width
(1) $w(\gamma):=\operatorname{arcsinh}\frac{1}{\sinh\frac{\ell(\gamma)}{2}}$
is always embedded in $X_{g}$. We call it the standard collar of $\gamma$. Moreover, $\operatorname{\mathcal{C}}(\gamma,w(\gamma))$ is isometric to $[-w(\gamma),w(\gamma)]\times\mathbb{R}/\mathbb{Z}$ endowed with the hyperbolic metric
$\mathop{}\!\mathrm{d}\rho^{2}+\ell(\gamma)^{2}\cosh^{2}\!\rho\mathop{}\!\mathrm{d}t^{2}.$
The pair of coordinates $(\rho,t)$ is called the Fermi coordinates on $\operatorname{\mathcal{C}}(\gamma,w(\gamma))$. For any $x\in X_{g}$, let $\operatorname{inj}(x)$ denote the injectivity radius at $x$. We recall that
Lemma 2.1 (Collar Lemma, [Bus92, Theorem 4.1.6]).
Let $\{\gamma_{1},\cdots,\gamma_{N}\}$ be the set of all simple closed geodesics of length $\leq 2\operatorname{arcsinh}1$ in $X_{g}$, and let $\operatorname{\mathcal{C}}_{i}$ be the standard collar of $\gamma_{i}$. Then the followings hold.
  1. (1)
    $N\leq 3g-3$.
  2. (2)
    The collars $\operatorname{\mathcal{C}}_{i}$ are pairwise disjoint.
  3. (3)
    Suppose $x\in\operatorname{\mathcal{C}}_{i}$ and $\operatorname{dist}(x,\partial\operatorname{\mathcal{C}}_{i})=d$, then
    (2) $\sinh\operatorname{inj}(x)=\cosh\frac{\ell(\gamma_{i})}{2}\cosh d-\sinh d.$
Without loss of generality, we now fix $\epsilon$ so that
(3) $0<\epsilon<\operatorname{arcsinh}\frac{1}{\sinh 2}\approx 0.272$
Recall that
$\mathcal{N}_{\epsilon}(X_{g})=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\}.$
Since $\epsilon<\operatorname{arcsinh}1$, by Part (2) of Lemma 2.1, $\mathcal{N}_{\epsilon}(X_{g})$ consists of pairwise disjoint simple closed geodesics. Moreover, by (1), for any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$ we have
$w(\gamma)>1.$
In the proof of Theorem 1.1, we will use a thick-thin decomposition of $X_{g}$ which is defined as follows.
Definition 2.2.
The thin part of $X_{g}$ is defined as
$X^{\mathrm{thin}}_{g}:=\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1).$
The thick part of $X_{g}$ is defined as the closure of its complement, which is given by
$X^{\mathrm{thick}}_{g}:=\overline{X\setminus X^{\mathrm{thin}}_{g}}=\mathcal{A}_{1}\sqcup\mathcal{A}_{2}\sqcup\cdots\sqcup\mathcal{A}_{I}.$
Here $I\geq 1$, and $\mathcal{A}_{k}$, $1\leq k\leq I$, are all components of $X^{\mathrm{thick}}_{g}$.
Then we have the following.
  • By (2) and (3), for any $x\in X^{\mathrm{thick}}_{g}$ we have
    (4) $\operatorname{inj}(x)\geq\min\big{\{}\epsilon,\,\operatorname{arcsinh}(\cosh 1-\sinh 1)\big{\}}=\epsilon.$
  • $\operatorname{\forall}\gamma\in\mathcal{N}_{\varepsilon}(X_{g})$, the length of a component of $\partial\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)$ is equal to
    (5) $\ell(\gamma)\cosh(w(\gamma)-1)\leq\frac{2}{e}\ell(\gamma)\cosh w(\gamma)$
    $=\frac{2}{e}\sqrt{\ell(\gamma)^{2}+\frac{\ell(\gamma)^{2}}{\sinh^{2}(\ell(\gamma)/2)}}\leq\frac{2}{e}\sqrt{1+4}<2.$
3. Mass distribution of eigenfunctions
One key ingredient in the proof of Theorem 1.1 is that, if $m(\lambda)$ is large (compared to the number of short closed geodesics $N_{\epsilon}$), then we can find a subspace of the eigenspace $E(\lambda)$ such that the $L^{2}$-norm of any eigenfunction in this subspace is concentrated on the thick part of $X_{g}$.
Let $f$ be an eigenfunction of $\Delta$ with eigenvalue $\lambda$, and let $\operatorname{\mathcal{C}}(\gamma,w)$ be an embedded collar in $X_{g}$. In the Fermi coordinates $(\rho,t)$, the Laplacian is given by
$-\Delta=\frac{\partial^{2}}{\partial\rho^{2}}+\tanh\rho\frac{\partial}{\partial\rho}+\frac{1}{\ell(\gamma)^{2}\cosh^{2}\!\rho}\frac{\partial^{2}}{\partial t^{2}}.$
Suppose $f$ has the following Fourier expansion on $\operatorname{\mathcal{C}}(\gamma,w)$,
(6) $f(\rho,t)=\alpha_{0}(\rho)+\sum_{j=1}^{\infty}\big{(}\alpha_{j}(\rho)\cos(2\pi jt)+\beta_{j}(\rho)\sin(2\pi jt)\big{)}.$
Then $\sqrt{\cosh\rho}\cdot\alpha_{j}$ and $\sqrt{\cosh\rho}\cdot\beta_{j}$ satisfy the following differential equation (one may refer to e.g., [Gam02, Section 4], [Mon15, Section 3] and [WZ25, Section 5] for more details):
(7) $\frac{\mathop{}\!\mathrm{d}^{2}u}{\mathop{}\!\mathrm{d}\rho^{2}}=\bigg{(}\frac{1}{4}-\lambda+\Big{(}\frac{1}{4}+\frac{4\pi^{2}j^{2}}{\ell(\gamma)^{2}}\Big{)}\frac{1}{\cosh^{2}\rho}\bigg{)}u.$
Let $\varphi_{j}$, $\psi_{j}$ be two linearly independent solutions of (7) satisfying
$\varphi_{j}(0)=\frac{\mathop{}\!\mathrm{d}\psi_{j}}{\mathop{}\!\mathrm{d}\rho}(0)=0,\textrm{\quad and\quad}\frac{\mathop{}\!\mathrm{d}\varphi_{j}}{\mathop{}\!\mathrm{d}\rho}(0)=\psi_{j}(0)=1.$
Recall that $w(\gamma)$ is the width of the standard collar of $\gamma$. We have the following elementary lemma:
Lemma 3.1.
Suppose that $\ell(\gamma)\leq 1$. If $j\geq\lfloor\sqrt{\lambda}\rfloor+1$, then
$\max\bigg{\{}\frac{\int_{0}^{w(\gamma)-1}\varphi_{j}^{2}\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\varphi_{j}^{2}\mathop{}\!\mathrm{d}\rho},\ \frac{\int_{0}^{w(\gamma)-1}\psi_{j}^{2}\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\psi_{j}^{2}\mathop{}\!\mathrm{d}\rho}\bigg{\}}\leq\frac{4}{e^{2}}.$
Proof.
Suppose that $\rho\leq w(\gamma)$. By (1) and our assumption $\ell(\gamma)\leq 1$, we have
$\ell(\gamma)^{2}\cosh^{2}\!\rho\leq\ell(\gamma)^{2}+\frac{\ell(\gamma)^{2}}{\sinh^{2}\!\frac{\ell(\gamma)}{2}}\leq 5.$
Therefore if $j\geq\lfloor\sqrt{\lambda}\rfloor+1$, then
$\frac{1}{4}-\lambda+\Big{(}\frac{1}{4}+\frac{4\pi^{2}j^{2}}{\ell(\gamma)^{2}}\Big{)}\frac{1}{\cosh^{2}\rho}\geq\frac{4\pi^{2}j^{2}}{5}-\lambda\geq 1.$
It follows that $\varphi_{j}$ and $\psi_{j}$ satisfy
$\frac{\mathop{}\!\mathrm{d}^{2}\varphi_{j}}{\mathop{}\!\mathrm{d}\rho^{2}}\geq\varphi_{j}\text{\quad and\quad}\frac{\mathop{}\!\mathrm{d}^{2}\psi_{j}}{\mathop{}\!\mathrm{d}\rho^{2}}\geq\psi_{j}.$
By plugging $u_{1}=\varphi_{j}\textit{ or }\psi_{j}$ and $u_{2}=\cosh\rho$ into [WZ25, Lemma 5.2], then we have
(8) $\max\bigg{\{}\frac{\int_{0}^{w(\gamma)-1}\varphi_{j}^{2}\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\varphi_{j}^{2}\mathop{}\!\mathrm{d}\rho},\ \frac{\int_{0}^{w(\gamma)-1}\psi_{j}^{2}\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\psi_{j}^{2}\mathop{}\!\mathrm{d}\rho}\bigg{\}}$ $\leq\frac{\int_{0}^{w(\gamma)-1}\cosh^{2}\!\rho\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\cosh^{2}\!\rho\mathop{}\!\mathrm{d}\rho}$
$=\frac{2w(\gamma)-2+\sinh(2w(\gamma)-2)}{2w(\gamma)+\sinh(2w(\gamma))}$
$\leq\frac{3\exp(2w(\gamma)-2)}{\exp(2w(\gamma))-\exp(-2w(\gamma))}$
$\leq\frac{4}{e^{2}},$
where we use the fact that $\exp(2w(\gamma))-\exp(-2w(\gamma))\geq 3\exp(2w(\gamma))/4$. The proof is complete. ∎
Next, for any $j\geq 0$ and any simple closed geodesic $\gamma\subset X_{g}$, we define
$\varphi_{j,\gamma}^{1}(x)=\begin{cases}\varphi_{j}(\rho)\cos(2\pi jt),&x=(\rho,t)\in\operatorname{\mathcal{C}}(\gamma,w(\gamma));\\ 0,&x\notin\operatorname{\mathcal{C}}(\gamma,w(\gamma)).\end{cases}$
For any $j\geq 1$, we define
$\varphi_{j,\gamma}^{2}(x)=\begin{cases}\varphi_{j}(\rho)\sin(2\pi jt),&x=(\rho,t)\in\operatorname{\mathcal{C}}(\gamma,w(\gamma));\\ 0,&x\notin\operatorname{\mathcal{C}}(\gamma,w(\gamma)).\end{cases}$
And let $\psi_{j,\gamma}^{1},\psi_{j,\gamma}^{2}\in L^{2}(X_{g})$ be defined similarly. For all $j\geq 0$, we now define a subspace $S(j,\gamma)$ of $L^{2}(X_{g})$ as
$S(j,\gamma):=\mathrm{span}\big{\{}\varphi_{0,\gamma}^{1},\varphi_{1,\gamma}^{1},\varphi_{1,\gamma}^{2},\cdots,\varphi_{j,\gamma}^{1},\varphi_{j,\gamma}^{2},\psi_{0,\gamma}^{1},\psi_{1,\gamma}^{1},\psi_{1,\gamma}^{2},\cdots,\psi_{j,\gamma}^{1},\psi_{j,\gamma}^{2}\big{\}}.$
Then we have
Lemma 3.2.
Let $f$ be an eigenfunction with eigenvalue $\lambda$, and let $\gamma$ be a simple closed geodesic of length $\leq 1$. If $f\in S(\lfloor\sqrt{\lambda}\rfloor,\gamma)^{\bot}$, the orthogonal complement of $S(\lfloor\sqrt{\lambda}\rfloor,\gamma)$, then
$\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A\leq 2\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma))\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A.$
Proof.
We use the same notation as in (6). Assume that
$\sqrt{\cosh\rho}\cdot\alpha_{j}(\rho)$ $=a_{j1}\cdot\varphi_{j}(\rho)+a_{j2}\cdot\psi_{j}(\rho),$
$\sqrt{\cosh\rho}\cdot\beta_{j}(\rho)$ $=b_{j1}\cdot\varphi_{j}(\rho)+b_{j2}\cdot\psi_{j}(\rho).$
Then the $L^{2}$-norm of $f$ on the collar $\operatorname{\mathcal{C}}(\gamma,w)$ is
(9) $\int_{\operatorname{\mathcal{C}}(\gamma,w)}f^{2}\mathop{}\!\mathrm{d}A=2\ell^{2}(\gamma)\cdot\int_{0}^{w}\Big{(}a_{01}^{2}\varphi_{0}^{2}+a_{02}^{2}\psi_{0}^{2}+\frac{1}{2}\sum_{j=1}^{\infty}(c_{j1}^{2}\varphi_{j}^{2}+c_{j2}^{2}\psi_{j}^{2})\Big{)}\mathop{}\!\mathrm{d}\rho,$
where $c_{ji}^{2}=a_{ji}^{2}+b_{ji}^{2}$. If $f\in S(\lfloor\sqrt{\lambda}\rfloor,\gamma)^{\bot}$, then
(10) $a_{01}=a_{02}=0,\text{\quad and\quad}c_{j1}=c_{j2}=0,\ 1\leq j\leq\lfloor\sqrt{\lambda}\rfloor.$
It then follows from (9), (10) and Lemma 3.1 that
$\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A$ $\leq\frac{4/e^{2}}{1-4/e^{2}}\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma))\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A$
$<2\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma))\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A.$
The proof is complete. ∎
Recall that for $10<\epsilon<$ 1,
$\mathcal{N}_{\epsilon}(X_{g})=\{\gamma;\ \text{$\gamma\subset X_{g}$ is a simple closed geodesic of length $<2\epsilon$}\}.$
We have the following lemma.
Lemma 3.3.
Let $X_{g}$ be a closed hyperbolic surface, and let $E(\lambda)\subset L^{2}(X_{g})$ be an eigenspace of $\Delta$ with respect to the eigenvalue $\lambda$. Then there exists a subspace $S$ of $E(\lambda)$ satisfying
$\operatorname{codim}S\leq 4(\lfloor\sqrt{\lambda}\rfloor+1)\cdot N_{\epsilon}(X_{g})$
such that $\operatorname{\forall}f\in S$ (if $S\neq 0$),
$\int_{X_{g}}f^{2}\mathop{}\!\mathrm{d}A\leq 3\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A,$
where $X^{\mathrm{thick}}_{g}$ is the thick part of $X_{g}$ corresponding to $\mathcal{N}_{\epsilon}(X_{g})$.
Proof.
Consider the orthogonal projection
$\Pi:E(\lambda)\to\bigoplus_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}S(\lfloor\sqrt{\lambda}\rfloor,\gamma).$
Then
$\operatorname{codim}(\ker\Pi)\leq\dim\Big{(}\bigoplus_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}S(\lfloor\sqrt{\lambda}\rfloor,\gamma)\Big{)}\leq 4(\lfloor\sqrt{\lambda}\rfloor+1)\cdot N_{\epsilon}(X_{g}),$
and by Lemma 3.2, for any $f\in\ker\Pi$ (if $\ker\Pi\neq 0$) we have
$\int_{X_{g}}f^{2}\mathop{}\!\mathrm{d}A$ $=\int_{\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A+\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A$
$\leq 2\int_{\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma))\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)}f^{2}\mathop{}\!\mathrm{d}A+\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A$
$\leq 3\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A.$
Thus, the subspace $\ker\Pi\subset E(\lambda)$ satisfies the desired properties. ∎
4. Proof of Theorem 1.1
In this section, we prove Theorem 1.1.
Let $X_{g}$ be a closed hyperbolic surface of genus $g$. Recall that
$\epsilon\in\Big{(}0,\operatorname{arcsinh}\frac{1}{\sinh 2}\Big{)}$
is a fixed constant, and $N_{\epsilon}(X_{g})=\#\mathcal{N}_{\epsilon}(X_{g})$. By the general upper bound of the first nonzero eigenvalue (e.g., [Che75, Corollary 2.3] or [WZ25, Proposition 9.1]), one has that as $g\to\infty$,
$\lambda_{1}(X_{g})\leq\frac{1}{4}+o(1).$
Now we always assume
$\lambda_{1}(X_{g})<1$
throughout this section by taking $g$ sufficiently large.
4.1. Rescaling the metric
We first rescale the hyperbolic metric on $X_{g}$ by $\epsilon^{-2}$ to obtain a new Riemannian surface $X_{g,\epsilon}$. Then the 2-dimensional simply connected space form of curvature $-\epsilon^{2}$, denoted by $\mathbb{H}_{\epsilon}$, is the universal cover of $X_{g,\epsilon}$. Let $\operatorname{dist}_{\epsilon}(\cdot,\cdot)$ denote the distance function on $X_{g,\epsilon}$, and for any $x\in X_{g,\epsilon}$, let $B_{\epsilon}(x,r)$ denote the geodesic ball of radius $r>0$ centered at $x$.
Let $\Delta_{\epsilon}$ be the Laplacian on $X_{g,\epsilon}$. Then $\lambda$ is an eigenvalue of the Laplacian $\Delta$ on $X_{g}$ if and only if $\epsilon^{2}\lambda$ is an eigenvalue of $\Delta_{\epsilon}$. It follows that
$m(\lambda_{1}(X_{g,\epsilon}))=m(\lambda_{1}(X_{g})).$
Hence, estimating the upper bound of $m(\lambda_{1}(X_{g}))$ is equivalent to estimating that of $m(\lambda_{1}(X_{g,\epsilon}))$. In the remainder of this section, we will work with $X_{g,\epsilon}$. Moreover, we use the notations
$X^{\mathrm{thick}}_{g,\epsilon},\ X^{\mathrm{thin}}_{g,\epsilon},\ \mathcal{A}_{k,\epsilon},\ \text{and\ }\operatorname{\mathcal{C}}(\gamma,w)_{\epsilon}$
to denote the images of
$X^{\mathrm{thick}}_{g},\ X^{\mathrm{thin}}_{g},\ \mathcal{A}_{k},\ \text{and\ }\operatorname{\mathcal{C}}(\gamma,w)$
under the rescaling by $\epsilon^{-2}$, respectively. For any $x\in X_{g,\epsilon}$, the Gaussian curvature
(11) $\mathrm{cur}(x)=-\epsilon^{2}>-1,$
and by (4), for any $x\in X^{\mathrm{thick}}_{g,\epsilon}$,
(12) $\operatorname{inj}(x)\geq 1.$
Let $r>0$. We recall the notions of $r$-net and $r$-separated set.
Definition 4.1.
Let $Y\subset X_{g,\epsilon}$ be a subset.
  1. (1)
    A set of points $\{x_{1},\cdots,x_{l}\}$ is called an $r$-net of $Y$ if for any $y\in Y$, there exists $x_{i}$ such that $\operatorname{dist}_{\epsilon}(x_{i},y)\leq r$.
  2. (2)
    A set of points $\{x_{1},\cdots,x_{l}\}$ is called an $r$-separated set if $\operatorname{dist}_{\epsilon}(x_{i},x_{j})\geq r$ for any $i\neq j$.
  3. (3)
    A set of points $\{x_{1},\cdots,x_{l}\}$ is called an $r$-separated net of $Y$ if it is both $r$-separated and an $r$-net of $Y$.
Let $\mathcal{A}_{k,\epsilon}$, $1\leq k\leq I$, be the components of $X^{\mathrm{thick}}_{g,\varepsilon}$. For all $r>0$ and $k\in[1,I]$, set
(13) $\mathcal{A}_{k,\epsilon}^{r}=\{x\in\mathcal{A}_{k,\epsilon}|\ \operatorname{dist}_{\epsilon}(x,X^{\mathrm{thin}}_{g,\epsilon})\geq r\}.$
Note that $\mathcal{A}_{k,\epsilon}^{r}$ might not be connected.
Lemma 4.2.
With the same notation as above, if $r\geq 4$ and $\mathcal{A}_{k,\epsilon}^{r}\neq\emptyset$, then $\mathcal{A}_{k,\epsilon}^{r}$ admits an $r$-separated net of cardinality at most
$\max\bigg{\{}1,\,\frac{16\operatorname{Area}(\mathcal{A}_{k,\epsilon})}{\pi\cdot r}\bigg{\}}.$
Proof.
The proof is standard. Firstly, if there exists $x\in\mathcal{A}_{k,\epsilon}^{r}$ such that $\mathcal{A}_{k,\epsilon}^{r}\subset B_{\epsilon}(x,r)$, then the singleton set $\{x\}$ is an $r$-net of $\mathcal{A}_{k,\epsilon}^{r}$. Otherwise, let
$\{x_{1},\cdots,x_{l}\}\subset\mathcal{A}_{k,\epsilon}^{r},\quad l\geq 2$
be a maximal $r$-separated set. This is clearly an $r$-separated net of $\mathcal{A}_{k,\epsilon}^{r}$. So we have
$B_{\epsilon}(x_{i},r/2)\cap B_{\epsilon}(x_{j},r/2)=\emptyset,\quad\operatorname{\forall}i\neq j.$
Moreover, by the definition of $\mathcal{A}_{k,\epsilon}^{r}$, we have $B_{\epsilon}(x_{i},r/2)\subset\mathcal{A}_{k,\epsilon}$, $\operatorname{\forall}1\leq i\leq l$, so
(14) $\sum_{i=1}^{l}\operatorname{Area}\big{(}B_{\epsilon}(x_{i},r/2)\big{)}\leq\operatorname{Area}(\mathcal{A}_{k,\epsilon}).$
Next, for each $1\leq i\leq l$, there exists $y_{i}\in\mathcal{A}_{k,\epsilon}$ such that $\operatorname{dist}_{\epsilon}(x_{i},y_{i})=r/2$. Let $\alpha_{i}:[0,r/2]\to X^{\mathrm{thick}}_{g,\epsilon}$ be a shortest geodesic segment from $x_{i}$ to $y_{i}$ parameterized by the arc length. Then for each integer $n\in\{0,\cdots,\lfloor r/2\rfloor-1\}$, we have
$B_{\epsilon}(\alpha_{i}(n),1/2)\subset B_{\epsilon}(x_{i},r/2)\subset\mathcal{A}_{k,\epsilon},$
and these geodesic balls are pairwise disjoint. By (11) and (12), $\mathrm{cur}(\alpha_{i}(n))=-\epsilon^{2}$ and $\operatorname{inj}(\alpha_{i}(n))\geq 1$, which implies
$\operatorname{Area}\big{(}B_{\epsilon}(\alpha_{i}(n),1/2)\big{)}=\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(1/2)\big{)}=\frac{2\pi}{\epsilon^{2}}(\cosh\frac{\epsilon}{2}-1)\geq\frac{\pi}{4},$
where $B_{\mathbb{H}_{\epsilon}}(1/2)$ denotes the geodesic ball of radius $1/2$ in $\mathbb{H}_{\epsilon}$. The last inequality follows from the estimate $\cosh\epsilon/2-1\geq\epsilon^{2}/8$. Thus,
$\operatorname{Area}\big{(}B_{\epsilon}(x_{i},r/2)\big{)}$ $\geq\sum_{n=0}^{\lfloor r/2\rfloor-1}\operatorname{Area}\big{(}B_{\epsilon}(\alpha_{i}(n),1/2)\big{)}$
$\geq\frac{\pi}{4}\cdot\Big{\lfloor}\frac{r}{2}\Big{\rfloor}\geq\frac{\pi}{16}\cdot r.$
Combining this with (14), we obtain the desired bound. ∎
Write $X_{g,\epsilon}=\mathbb{H}_{\epsilon}/\Gamma$, and let $K_{t,\epsilon}(\cdot,\cdot)$ denote the heat kernel on $X_{g,\epsilon}$. Then for any $f\in L^{2}(X_{g,\epsilon})$,
$(e^{-t\Delta_{\epsilon}}f)(x)=\int_{X_{g,\epsilon}}K_{t,\epsilon}(x,y)f(y)\mathop{}\!\mathrm{d}y.$
Moreover, let $k_{t,\epsilon}(\cdot,\cdot)$ denote the heat kernel on $\mathbb{H}_{\epsilon}$, then
(15) $K_{t,\epsilon}(x,y)=\sum_{\gamma\in\Gamma}k_{t,\epsilon}(\gamma\tilde{x},\tilde{y}),$
where $\tilde{x}$ and $\tilde{y}$ are lifts of $x$ and $y$, respectively. Since $k_{t,\epsilon}(\tilde{x},\tilde{y})$ depends only on $\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})$, we also write it as $k_{t,\epsilon}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}))$. Next, we present a standard estimate for the heat kernel on $X_{g,\epsilon}$ which will be used in the subsequent analysis.
Lemma 4.3.
There exists a universal constant $C_{1}>0$, independent of $\epsilon$, such that for any point $x\in X_{g,\epsilon}$ with $\operatorname{inj}(x)\geq 1$ and any $t\geq 1$,
$\|K_{t,\epsilon}(x,\cdot)\|_{L^{\infty}(X_{g,\epsilon})}\leq C_{1}\exp(4t).$
Proof.
Since
$k_{t,\epsilon}(\tilde{x},\tilde{y})=k_{t,\epsilon}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y}))=\epsilon^{2}k_{\epsilon^{2}t,1}(\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})),$
it follows from [Ber16, Lemma 3.13] that for all $t\geq 1$,
(16) $k_{t,\epsilon}(\tilde{x},\tilde{y})\ll\exp(-\frac{\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})^{2}}{8\epsilon^{2}t})\ll\exp(-\frac{\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\tilde{x},\tilde{y})^{2}}{8t}),\quad\operatorname{\forall}\tilde{x},\tilde{y}\in\mathbb{H}_{\epsilon}.$
For any $m\geq 0\in\mathbb{N}$ and any $\tilde{x},\tilde{y}\in\mathbb{H}_{\epsilon}$, since $\operatorname{inj}(x)\geq 1$,
(17) #$\#\{\gamma\in\Gamma|\ m\leq\operatorname{dist}_{\mathbb{H}_{\epsilon}}(\gamma\tilde{x},\tilde{y})<m+1\}\leq\frac{\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(m+1+1/2)\big{)}}{\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(1/2)\big{)}}$
$\ll\operatorname{Area}\big{(}B_{\mathbb{H}_{\epsilon}}(m+1+1/2)\big{)}\ll\operatorname{Area}\big{(}B_{\mathbb{H}}(m+1+1/2)\big{)}\ll\exp(m).$
Therefore, by (15), (16) and (17),
$K_{t,\epsilon}(x,y)$ $\ll\sum_{m=0}^{\infty}\exp(m-\frac{m^{2}}{8t})$
$\leq(4t+1)\exp(2t)+\sum_{m\geq 4t}\exp(8t-m)$
$\ll\exp(4t).$
The proof is complete. ∎
For simplicity, we write $\lambda_{1,\epsilon}$ for $\lambda_{1}(X_{g,\epsilon})=\epsilon^{2}\lambda_{1}(X_{g})$, and omit $X_{g}$ in $\lambda_{1}(X_{g})$, $\mathcal{N}_{\epsilon}(X_{g})$, and $N_{\epsilon}(X_{g})$ for the remainder of this section. Set
(18) $r_{1}=c\log\log\Big(\frac{Kg}{N_{\epsilon}+1}\Big{missing}),\qquad r_{2}=c\log\Big(\frac{Kg}{N_{\epsilon}+1}\Big{missing}),$
where $c>0$ is a universal constant that will be chosen later. We assume that $K\geq 10$ is a sufficiently large universal constant such that $\frac{Kg}{N_{\epsilon}+1}$ is large enough for our analysis, particularly to ensure that $r_{1}\gg 1$ and $r_{2}/r_{1}\gg 1$.
4.2. Reduce the problem to the thick part
In this subsection, we show that it suffices to consider only the thick part of $X_{g,\epsilon}$. Set
$X_{g,\epsilon}^{r_{1}}:=\{x\in X^{\mathrm{thick}}_{g,\epsilon}|\ \operatorname{dist}_{\epsilon}(x,X^{\mathrm{thin}}_{g,\epsilon})\geq r_{1}\}.$
First, we have the following lemma.
Lemma 4.4.
There exists an $r_{1}$-net $\{x_{1},\cdots,x_{l}\}$ of $X_{g,\epsilon}^{r_{1}}$ such that
$l\ll\frac{1}{\epsilon^{2}}\frac{g}{\log\log\frac{10g}{N_{\epsilon}+1}}.$
Proof.
By definition we have
$X_{g,\epsilon}^{r_{1}}=\mathcal{A}_{1,\epsilon}^{r_{1}}\sqcup\mathcal{A}_{2,\epsilon}^{r_{1}}\sqcup\cdots\sqcup\mathcal{A}_{I,\epsilon}^{r_{1}},\quad I\geq 1,$
where $\mathcal{A}_{k,\epsilon}^{r_{1}}$ is defined by (13). Clearly, the union of $r_{1}$-nets of $\mathcal{A}_{k,\epsilon}^{r_{1}}$’s is an $r_{1}$-net of $X_{g,\epsilon}^{r_{1}}$. Thus by Lemma 4.2,
$l$ $\leq\sum_{k=1}^{I}\max\bigg{\{}1,\frac{16\operatorname{Area}(\mathcal{A}_{k,\epsilon})}{\pi\cdot r_{1}}\bigg{\}}\leq\sum_{k=1}^{I}\bigg{(}1+\frac{16\operatorname{Area}(\mathcal{A}_{k,\epsilon})}{\pi\cdot r_{1}}\bigg{)}$
$\leq N_{\epsilon}+1+\frac{64\pi\epsilon^{-2}(g-1)}{\pi\cdot r_{1}}\ll\frac{1}{\epsilon^{2}}\frac{g}{\log\log\frac{10g}{N_{\epsilon}+1}},$
where we use the bounds
$I\leq N_{\epsilon}+1\text{\quad and\quad}\sum_{k=1}^{I}\operatorname{Area}(\mathcal{A}_{k,\epsilon})\leq\operatorname{Area}(X_{g,\epsilon})=4\pi\epsilon^{-2}(g-1).$
The proof is complete. ∎
Next, we extend $\{x_{1},\cdots,x_{l}\}$ to a maximal 1-separated set of $X_{g,\epsilon}$, denoted by
$\{x_{1},\cdots,x_{l},x_{l+1},\cdots,x_{s}\}$
where $l<s\in\mathbb{Z}^{+}$. For each $i\in\{1,\cdots,s\}$, define
$V_{i}:=\big{\{}y\in X_{g,\epsilon}|\ \operatorname{dist}_{\epsilon}(y,x_{i})\leq\operatorname{dist}_{\epsilon}(y,x_{j}),\ \forall j\in\{1,\cdots,s\}\big{\}}.$
Since $\{x_{1},\cdots,x_{s}\}$ is a 1-separated set,
$B_{\epsilon}(x_{i},1/2)\subset V_{i}\subset B_{\epsilon}(x_{i},1),\quad\forall i\in\{1,\cdots,s\}.$
For each $i\in\{1,\cdots,l\}$, set
$\psi_{i}=\frac{1}{\sqrt{\operatorname{Area}(V_{i})}}\textbf{1}_{\mathring{V}_{i}},$
where $\textbf{1}_{\mathring{V}_{i}}$ denotes the indicator function of the interior $\mathring{V}_{i}$ of $V_{i}$. Then $\norm{\psi_{i}}_{L^{2}}=1$. Moreover, $\operatorname{\forall}1\leq i\leq l$, since $\operatorname{inj}(x_{i})\geq 1$, it follows that
(19) $\operatorname{Area}(V_{i})\geq\operatorname{Area}\big{(}B_{\epsilon}(x_{i},1/2)\big{)}=\frac{2\pi}{\epsilon^{2}}(\cosh\frac{\epsilon}{2}-1)\geq\frac{\pi}{4}.$
Let $P$ be the orthogonal projection from $L^{2}(X_{g,\epsilon})$ to $\mathrm{span}\{\psi_{1},\cdots,\psi_{l}\}^{\bot}$, i.e.
$Pf=f-\sum_{k=1}^{l}\langle f,\psi_{k}\rangle\psi_{k},\quad\forall f\in L^{2}(X_{g,\epsilon}).$
Let $\delta_{x}$ denote the Dirac delta distribution at $x$, then $P\delta_{x}$ is given as
$P\delta_{x}=\delta_{x}-\sum_{k=1}^{l}\psi_{k}(x)\psi_{k}.$
Following the ideas of [LM24, JTY+21], we will consider the multiplicity of the eigenvalue $e^{-r_{1}\lambda_{1,\epsilon}}$ in the spectrum of the operator $Pe^{-r_{1}\Delta_{\epsilon}}P$. First, we have
Lemma 4.5.
If $m(\lambda_{1,\epsilon})>l+4N_{\epsilon}$, then there exists a subspace $S\subset E(\lambda_{1,\epsilon})$ with
$\dim S\geq m(\lambda_{1,\epsilon})-l-4N_{\epsilon},$
such that each $\varphi\in S$ is an eigenfunction of $Pe^{-r_{1}\Delta_{\epsilon}}P$ with eigenvalue $e^{-r_{1}\lambda_{1,\epsilon}}$, and
(20) $\int_{X_{g,\epsilon}}\varphi^{2}\mathrm{d}A\leq 3\int_{X^{\mathrm{thick}}_{g,\epsilon}}\varphi^{2}\mathrm{d}A.$
Proof.
By Lemma 3.3, there exists a subspace $S_{1}\subset E(\lambda_{1})$ with
$\dim S_{1}\geq m(\lambda_{1})-4(\lfloor\sqrt{\lambda}_{1}\rfloor+1)N_{\epsilon}=m(\lambda_{1})-4N_{\epsilon},$
such that every $\varphi\in S_{1}$ satisfies the inequality
$\int_{X_{g}}\varphi^{2}\mathrm{d}A\leq 3\int_{X^{\mathrm{thick}}_{g}}\varphi^{2}\mathrm{d}A.$
Since each $\varphi\in E(\lambda_{1})$ corresponds to an eigenfunction $\varphi_{\epsilon}\in E(\lambda_{1,\epsilon})$ under the rescaling by $\epsilon^{-2}$, we obtain a subspace $S_{1,\epsilon}\subset E(\lambda_{1,\epsilon})$ with
$\dim S_{1,\epsilon}=\dim S_{1}\geq m(\lambda_{1})-4N_{\epsilon},$
such that every function in $S_{1,\epsilon}$ satisfies the inequality (20).
Next, consider the orthogonal projection
$\Pi:S_{1,\epsilon}\to\mathrm{Span}\{\psi_{1},\cdots,\psi_{l}\}.$
Then
$\dim(\ker\Pi)\geq\dim S_{1,\epsilon}-l\geq m(\lambda_{1})-l-4N_{\epsilon},$
and every $\varphi\in\ker\Pi$ satisfies
$P\varphi=(\mathrm{Id}-\Pi)\varphi=\varphi,$
which implies
$Pe^{-r_{1}\Delta_{\epsilon}}P\varphi=Pe^{-r_{1}\Delta_{\epsilon}}\varphi=P(e^{-r_{1}\lambda_{1,\epsilon}}\varphi)=e^{-r_{1}\lambda_{1,\epsilon}}\varphi.$
Therefore, the subspace $\ker\Pi$ satisfies the desired properties. ∎
Let $\{\varphi_{j}\}_{j=1}^{\infty}$ be an orthonormal basis of $L^{2}(X_{g,\epsilon})$ consisting of eigenfunctions of the self-adjoint compact operator $Pe^{-r_{1,\epsilon}\Delta}P$, with the associated eigenvalues $\{\Lambda_{j}\}_{j=1}^{\infty}$. According to Lemma 4.5, one may choose this basis such that
(21) $\mathrm{Span}\{\varphi_{1},\cdots,\varphi_{\dim S}\}=S\text{\quad and\quad}\Lambda_{1}=\cdots=\Lambda_{\dim S}=e^{-r_{1}\lambda_{1,\epsilon}}.$
To give an upper bound of $m$, it suffices to give an upper bound of $\dim S$. Set
$m^{\prime}=\dim S\text{\quad and\quad}n=\lfloor r_{2}/r_{1}\rfloor.$
We have the following two lemmas.
Lemma 4.6.
Let $\delta_{x}$ denote the Dirac delta distribution at $x$. Then
$\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}=\sum_{j=1}^{\infty}\Lambda_{j}^{2n+2}\varphi_{j}^{2}.$
Proof.
The claim is based on a direct computation. For any $f\in L^{2}(X_{g,\epsilon})$, suppose that
$f=\sum_{j=1}^{\infty}a_{j}\varphi_{j},\quad a_{j}\in\mathbb{R},$
then since $Pe^{-r_{1}\Delta_{\epsilon}}P$ is self-adjoint,
$\big{\langle}(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x},\ f\big{\rangle}=((Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}f)(x)$
$=\sum_{j=1}^{\infty}a_{j}\Lambda_{j}^{n+1}\varphi_{j}(x)=\Big{\langle}\sum_{j=1}^{\infty}\Lambda_{j}^{n+1}\varphi_{j}(x)\varphi_{j},\ f\Big{\rangle}.$
Thus, we have
$\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}=\Big{\|}\sum_{j=1}^{\infty}\Lambda_{j}^{n+1}\varphi_{j}(x)\varphi_{j}\Big{\|}^{2}_{L^{2}(X_{g,\epsilon})}=\sum_{j=1}^{\infty}\Lambda_{j}^{2n+2}\varphi_{j}^{2},$
as desired. ∎
Lemma 4.7.
With the same notation as above,
$m^{\prime}\cdot e^{-r_{1}\lambda_{1,\epsilon}(2n+2)}\leq 3\int_{X^{\mathrm{thick}}_{g,\epsilon}}\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}\mathrm{d}A_{x}.$
Proof.
By Lemma 4.5 and Lemma 4.6, we have
$m^{\prime}\cdot e^{-r_{1}\lambda_{1,\epsilon}(2n+2)}$ $=e^{-r_{1}\lambda_{1,\epsilon}(2n+2)}\int_{X_{g,\epsilon}}\sum_{j=1}^{m^{\prime}}\varphi_{j}^{2}\mathop{}\!\mathrm{d}A$
[…]
$=3\int_{X^{\mathrm{thick}}_{g,\epsilon}}\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}\mathop{}\!\mathrm{d}A_{x}.$
The proof is complete. ∎
4.3. Estimate of the integral on the thick part
In this subsection, we give an upper bound for
$\int_{X^{\mathrm{thick}}_{g,\epsilon}}\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}\mathrm{d}A_{x}.$
We decompose $X_{g,\epsilon}^{\mathrm{thick}}$ into the following two disjoint sets (see Figure 3):
  • $Y_{1}:=\{x\in X^{\mathrm{thick}}_{g,\epsilon}|\ \operatorname{dist}_{\epsilon}(x,X^{\mathrm{thin}}_{g,\epsilon})\leq r_{1}+C^{\prime}r_{2}+2\}$;
  • $Y_{2}:=X_{g,\epsilon}^{\mathrm{thick}}\setminus Y_{1}$,
where $C^{\prime}>0$ is a universal constant that will be fixed later.
Figure 3. An illustration for $Y_{1}$ and $Y_{2}$
4.3.1. Estimate on $Y_{1}$
We first give an upper bound for
$\int_{Y_{1}}\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}^{2}_{L^{2}(X_{g,\epsilon})}\mathrm{d}A_{x}.$
We have the following two lemmas.
Lemma 4.8.
For any $x\in Y_{1}$,
$\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}_{L^{2}(X_{g,\epsilon})}\ll 1.$
Proof.
Since the operator norms of $P$, $e^{-r_{1}\Delta_{\epsilon}}$ and $e^{-(r_{1}-1)\Delta_{\epsilon}}$ are all no more than $1$, we have
(22) $\|(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}\|_{L^{2}(X_{g,\epsilon})}\leq\|e^{-\Delta_{\epsilon}}P\delta_{x}\|_{L^{2}(X_{g,\epsilon})}.$
Let $k_{x}\in\{1,\cdots,l\}$ be the unique index such that $x\in\mathring{V}_{k_{x}}$, if such an index exists. In this case, we have
(23) $(e^{-\Delta_{\epsilon}}P\delta_{x})(y)$ $=e^{-\Delta_{\epsilon}}\delta_{x}-\psi_{k_{x}}(x)\cdot(e^{-\Delta_{\epsilon}}\psi_{k_{x}})(y)$
$=K_{1,\epsilon}(x,y)-\psi_{k_{x}}(x)\int_{X_{g,\epsilon}}K_{1,\epsilon}(y,z)\psi_{k_{x}}(z)\mathrm{d}A_{z}.$
Then by (19), (23) and Lemma 4.3, we have
(24) $\|e^{-\Delta_{\epsilon}}P\delta_{x}\|_{L^{2}(X_{g,\epsilon})}$
[…]
If no such $k_{x}$ exists, then $e^{-\Delta_{\epsilon}}P\delta_{x}=e^{-\Delta_{\epsilon}}\delta_{x}$, and hence
(25) $\|e^{-\Delta_{\epsilon}}P\delta_{x}\|_{L^{2}(X_{g,\epsilon})}=\|K_{1,\epsilon}(x,\cdot)\|_{L^{2}(X_{g,\epsilon})}\leq\sqrt{C}_{1}.$
The lemma then follows from (22), (23) (24) and (25). ∎
Lemma 4.9.
Under the above assumptions,
$\int_{Y_{1}}\|(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}\|_{L^{2}(X_{g,\epsilon})}^{2}\mathrm{d}A_{x}\ll\frac{N_{\epsilon}\exp(2C^{\prime}r_{2}\epsilon)}{\epsilon^{2}}.$
Proof.
For each $\gamma\in\mathcal{N}_{\epsilon}$, let $x_{\gamma,\epsilon}$ and $y_{\gamma,\epsilon}$ be two points in distinct components of $\partial\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)_{\epsilon}$. Since by (5) and rescaling, the length of each component of $\partial\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1)_{\epsilon}$ is less than $2\epsilon^{-1}$, we have
$Y_{1}\subset\bigcup_{\gamma\in\mathcal{N}_{\epsilon}}\big{(}B_{\epsilon}(x_{\gamma,\epsilon},r_{1}+C^{\prime}r_{2}+2+\epsilon^{-1})\cup B_{\epsilon}(y_{\gamma,\epsilon},r_{1}+C^{\prime}r_{2}+2+\epsilon^{-1})\big{)}.$
Denote $R_{\epsilon}=r_{1}+C^{\prime}r_{2}+2+\epsilon^{-1}$, then,
(26) $\operatorname{Area}(Y_{1})$ $\leq\sum_{\gamma\in\mathcal{N}_{\epsilon}}\big{(}\operatorname{Area}(B_{\epsilon}(x_{\gamma,\epsilon},R_{\epsilon}))+\operatorname{Area}(B_{\epsilon}(y_{\gamma,\epsilon},R_{\epsilon}))\big{)}$
$\leq 2N_{\epsilon}\cdot\operatorname{Area}(B_{\mathbb{H}_{\epsilon}}(2C^{\prime}r_{2}+3\epsilon^{-1}))$
$=2N_{\epsilon}\cdot\frac{2\pi}{\epsilon^{2}}(\cosh(2C^{\prime}r_{2}\epsilon+3)-1)$
$\ll N_{\epsilon}\epsilon^{-2}\exp(2C^{\prime}r_{2}\epsilon),$
where $B_{\mathbb{H}_{\epsilon}}(r)$ denotes the geodesic ball of radius $r$ in $\mathbb{H}_{\epsilon}$. The conclusion then follows from (26) and Lemma 4.8. ∎
4.3.2. Estimate on $Y_{2}$
We omit the subscript $L^{2}(X_{g,\epsilon})$ in the following for simplicity. Following the approach in [LM24, Section 3 and Section 4] of Letrouit-Machado, we give an upper bound for
$\int_{Y_{2}}\|(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}\|^{2}\mathrm{d}A_{x}.$
For any $x\in Y_{2}$, let $\chi_{x}$ be the indicator function of the subset
$\bigcup_{1\leq i\leq s,V_{i}\cap B_{\epsilon}(x,C^{\prime}r_{2})\neq\emptyset}V_{i}.$
Note that
$\mathrm{supp}(\chi_{x})\subset B_{\epsilon}(x,C^{\prime}r_{2}+2),$
by the definition of $Y_{2}$, this implies that
$\mathrm{supp}(\chi_{x})\subset X_{g,\epsilon}^{r_{1}},\quad\forall x\in Y_{2}.$
Thus, every point $y\in\mathrm{supp}(\chi_{x})$ satisfies
$\operatorname{inj}(y)\geq 1.$
Recall the following facts:
  • Since $X_{g,\epsilon}^{r_{1}}\subset X^{\mathrm{thick}}_{g,\epsilon}$, it follows from (11) and (12) that for all $x\in X_{g,\epsilon}^{r_{1}}$, we have $\mathrm{cur}(x)>-1$ and $\operatorname{inj}(x)\geq 1$.
  • $\{x_{1},\cdots,x_{l}\}$ is an $r_{1}$-net of $X_{g,\epsilon}^{r_{1}}$, hence is also an $r_{1}$-net of $Y_{2}$;
  • $\|\psi_{i}\|_{L^{\infty}(X_{g})}\leq 2/\sqrt{\pi}$ for all $1\leq i\leq l$.
These facts allow us to repeat the arguments in the proofs of [LM24, Lemma 3.1, 3.4, 3.8], taking $\rho=1$ and $b=-1$ in the initial assumptions of [LM24, Section 3], to derive the following three special cases of those lemmas.
Lemma 4.10 ([LM24], Lemma 3.1).
There exists a universal constant $C_{2}>0$ such that if $C^{\prime}\geq 48$, then for any $x\in Y_{2}$,
$\norm{(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}}$
$\leq C_{2}\bigg{(}\sup_{\|\varphi\|=1}\norm{(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi}+\exp(-\frac{C^{\prime 2}r_{2}}{32})\bigg{)}.$
Following [LM24], we denote by $\varphi_{x}\in L^{2}(X_{g,\epsilon})$ a function that attains the following supremum
$\sup_{\|\varphi\|=1}\norm{P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P\varphi}.$
Lemma 4.11 ([LM24], Lemma 3.4).
Assume that $C^{\prime 2}\geq 80+48C^{\prime}$. Then there exists a universal constant $C_{3}>0$, and a subset $Y_{3}\subset Y_{2}$ with
$\operatorname{Area}(Y_{3})\ll\exp(3C^{\prime}r_{2}),$
such that for any $x\in Y_{2}\setminus Y_{3}$,
$\norm{e^{-r_{1}\Delta_{\epsilon}}|\varphi_{x}|}^{2}\leq e^{-2r_{1}\lambda_{1,\epsilon}}+C_{3}\exp(-C^{\prime}r_{2}).$
Lemma 4.12 ([LM24], Lemma 3.8).
There exist two universal constants $C_{4},C_{5}>0$ such that for any $C>16$ and any $x\in Y_{2}$,
$\|P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P\varphi_{x}\|^{2}$
$\leq\big{(}1-C_{4}\exp(-hr_{1})\big{)}\cdot\norm{e^{-r_{1}\Delta_{\epsilon}}|\varphi_{x}|}^{2}+C_{5}\exp(r_{1}-\frac{C^{2}r_{1}}{64}),$
where $h=50(1+C)$.
We now estimate the integral on $Y_{2}$. Assume that
(27) $C^{\prime}>48,\qquad C>16,$
(28) $c<\min\left(\frac{1}{8+6C^{\prime}},\frac{1}{4h}\right),$
(29) $C^{\prime 2}>\max\left(80+48C^{\prime},\frac{16}{c}\right),$
(30) $\frac{C^{2}}{64}>h+4,$
for example, $C^{\prime}=10^{5},C=10^{4},c=10^{-7}$. First, by Lemma 4.10 and Lemma 4.11,
(31) $\int_{Y_{2}}\|(Pe^{-r_{1}\Delta_{\epsilon}}P)^{n+1}\delta_{x}\|^{2}\mathrm{d}A_{x}$
[…]
where we use the bounds
$\operatorname{Area}(Y_{2})\ll\epsilon^{-2}g\text{\quad and\quad}\|(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}\|\leq 1,\ \operatorname{\forall}x\in Y_{2}.$
Next, we use Lemma 4.12 to give the following bound.
Lemma 4.13.
Under the above assumptions, if $1\ll r_{1}\ll r_{2}$, then
$\int_{Y_{2}\setminus Y_{3}}\|(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}\|^{2}\mathrm{d}A_{x}\ll\epsilon^{-2}g\cdot\exp(-\frac{1}{4}C_{4}\exp(-hr_{1})n)e^{-2r_{1}n\lambda_{1,\epsilon}}.$
Proof.
First, if $r_{2}/r_{1}$ is sufficiently large , then
$C_{3}\exp(-C^{\prime}r_{2})\leq\frac{1}{2}C_{4}\exp(-hr_{1})e^{-2r_{1}\lambda_{1,\epsilon}}.$
Thus by Lemma 4.11, for any $x\in Y_{2}\setminus Y_{3}$,
$\big{(}1-C_{4}\exp(-hr_{1})\big{)}\cdot\norm{e^{-r_{1}\Delta_{\epsilon}}|\varphi_{x}|}^{2}$
[…]
Next, by (30) and our assumption $\lambda_{1,\epsilon}=\epsilon^{2}\lambda_{1}<1$, if $r_{1}$ is sufficiently large, then
$C_{5}\exp(r_{1}-\frac{C^{2}r_{1}}{64})\leq C_{5}\exp\big(-hr_{1}-3r_{1}\big{missing})\leq\frac{1}{4}C_{4}\exp(-hr_{1})e^{-2r_{1}\lambda_{1,\epsilon}}.$
Combining the above two estimates, it follows from Lemma 4.12 that
$\|P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P\varphi_{x}\|^{2}\leq\Big{(}1-\frac{1}{4}C_{4}\exp(-hr_{1})\Big{)}\cdot e^{-2r_{1}\lambda_{1,\epsilon}}.$
Finally, by (28), we obtain
$\int_{Y_{2}\setminus Y_{3}}\|(P\chi_{x}e^{-r_{1}\Delta_{\epsilon}}\chi_{x}P)^{n}\varphi_{x}\|^{2}\mathrm{d}A_{x}$
[…]
as desired. ∎
Now we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
Combining Lemma 4.7, Lemma 4.9, (31) and Lemma 4.13, we have
(32) $m^{\prime}\cdot e^{-r_{1}\lambda_{1,\epsilon}(2n+2)}\ll{}$ $\epsilon^{-2}g\cdot\exp(-\frac{1}{4}C_{4}\exp(-hr_{1})n)e^{-2r_{1}n\lambda_{1,\epsilon}}$
$+\exp(3C^{\prime}r_{2})+\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16})+\epsilon^{-2}N_{\epsilon}\exp(2C^{\prime}r_{2}\epsilon)$
[…]
$+\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16})+\epsilon^{-2}(N_{\epsilon}+1)\exp(3C^{\prime}r_{2}).$
To obtain an upper bound for $m^{\prime}$, it suffices to control the following three terms:
  • $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\cdot\epsilon^{-2}(N_{\epsilon}+1)\exp(3C^{\prime}r_{2})$,
  • $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\cdot\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16})$,
  • $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\cdot\epsilon^{-2}g\exp(-\frac{1}{4}C_{4}\exp(-hr_{1})n)e^{-2r_{1}n\lambda_{1,\epsilon}}$.
Recall that $\lambda_{1,\epsilon}<1$ and $n=\lfloor r_{2}/r_{1}\rfloor$. It follows from (28) that
(33) $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\cdot\epsilon^{-2}(N_{\epsilon}+1)\exp(3C^{\prime}r_{2})\leq\epsilon^{-2}(N_{\epsilon}+1)\exp((4+3C^{\prime})r_{2})$
[…]
Similarly, by (28) and (29),
(34) $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\cdot\epsilon^{-2}g\exp(-\frac{C^{\prime 2}r_{2}}{16})\leq\epsilon^{-2}g\exp(\Big{(}4-\frac{C^{\prime 2}}{16}\Big{)}r_{2})$
[…]
For the third term, if $1\ll r_{1}\ll r_{2}$, then by (28),
(35) $e^{r_{1}\lambda_{1,\epsilon}(2n+2)}\epsilon^{-2}g\exp(-\frac{1}{4}C_{4}\exp(-hr_{1})n)e^{-2r_{1}n\lambda_{1,\epsilon}}$
[…]
where in the second inequality we apply
$\Big{(}\log\frac{Kg}{N_{\epsilon}+1}\Big{)}^{\frac{2}{3}}\ll n\exp(-\frac{1}{4c}r_{1})\text{\quad and\quad}2r_{1}\ll\frac{1}{20}C_{4}\Big{(}\log\frac{Kg}{N_{\epsilon}+1}\Big{)}^{\frac{2}{3}}.$
Finally, combining (32), (33), (34), (35), Lemma 4.4 and Lemma 4.5, we obtain
$m$ $\ll l+4N_{\epsilon}+\epsilon^{-2}(g(N_{\epsilon}+1))^{\frac{1}{2}}+\epsilon^{-2}\frac{g}{\log\frac{Kg}{N_{\epsilon}+1}}$
$\ll\epsilon^{-2}\frac{g}{\log\log\frac{10g}{N_{\epsilon}+1}}+N_{\epsilon}+\epsilon^{-2}(g(N_{\epsilon}+1))^{\frac{1}{2}}+\epsilon^{-2}\frac{g}{\log\frac{10g}{N_{\epsilon}+1}}$
$\ll\epsilon^{-2}\frac{g}{\log\log\frac{10g}{N_{\epsilon}+1}}.$
This completes the proof of Theorem 1.1. ∎
5. Proof of Theorem 1.2
In this section, we prove Theorem 1.2.
We use a thick-thin decomposition different from that used in the proof of Theorem 1.1. For any $\delta\in(0,1/2)$, let
$\epsilon(\delta)=\operatorname{arcsinh}\frac{1}{\sinh(1/\delta+2)}.$
Then for any $\epsilon\in(0,\epsilon(\delta))$, we define
(36) $X^{\mathrm{thin}}_{g}$ $:=\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-\frac{1}{\delta}),$
$X^{\mathrm{thick}}_{g}$ $:=\overline{X\setminus X^{\mathrm{thin}}_{g}}=\mathcal{A}_{1}\sqcup\mathcal{A}_{2}\sqcup\cdots\sqcup\mathcal{A}_{I_{\epsilon}},\quad I_{\epsilon}\geq 1.$
Next, for each $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, we define shell of $\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1/\delta)$ as
$\mathcal{S}(\gamma):=\operatorname{\mathcal{C}}(\gamma,w(\gamma)-\frac{1}{\delta})\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-\frac{1}{\delta}-2).$
Let $\mathcal{S}_{k}$ denote the union of the connected components of the shells adjacent to $\mathcal{A}_{k}$.
Since $\mathcal{A}_{k}$ is connected, for any $p,q\in\mathcal{A}_{k}$, there exists a shortest geodesic segment joining them that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. We have the following lemma:
Lemma 5.1.
Suppose $p,q\in\mathcal{A}_{k}$, and let $\alpha$ be a shortest geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. Then for any $x\in\alpha$,
$\operatorname{inj}(x)>\epsilon\text{\quad and\quad}B(x,\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}.$
Proof.
For any $x\in\alpha$, if $x\notin\operatorname{\mathcal{C}}(\gamma,w(\gamma))$ for all $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, then the lemma follows by definition. If $x\in\operatorname{\mathcal{C}}(\gamma,w(\gamma))$ for some $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, by considering the restriction of $\alpha$ to $X^{\mathrm{thin}}_{g}$, one may assume that both $p$ and $q$ are in one component of the boundary $\partial\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1/\delta)$. See Figure 4 for an illustration.
Figure 4. An illustration for the position of $\alpha$.
We first claim that
(37) $d:=\operatorname{dist}\big{(}x,\partial\operatorname{\mathcal{C}}(\gamma,w(\gamma))\big{)}\leq 1/\delta+1.$
Otherwise, $\ell(\alpha)\geq 2$, while by (5), $\ell(\alpha_{pq})<2$, which leads to a contradiction since we assume that $\alpha$ is shortest. Thus, $d\leq 1/\delta+1$, and by (2) we have
$\operatorname{inj}(x)$ $=\operatorname{arcsinh}\Big{(}\cosh\frac{\ell(\gamma)}{2}\cosh d-\sinh d\Big{)}$
$\geq\operatorname{arcsinh}(\cosh d-\sinh d)$
$\geq\operatorname{arcsinh}\frac{1}{e^{1/\delta+1}}$
$>\operatorname{arcsinh}\frac{1}{\sinh(1/\delta+2)}=\epsilon(\delta)>\epsilon.$
Moreover, since $\epsilon<1$, by (37) we have $B(x,\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}$, this concludes the proof of the lemma. ∎
Let $f$ be a normalized eigenfunction of $\Delta$ with eigenvalue $\lambda<1/4$, then we have the following lemma, which estimates the oscillations of $f$ over the components of the thick part of $X_{g}$.
Lemma 5.2.
Suppose $p,q\in\mathcal{A}_{k}$, and let $\alpha$ be a shortest geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. Then
|$\absolutevalue{f(p)-f(q)}^{2}\leq c(\epsilon)\cdot\lambda\cdot\ell(\alpha)$
where $c(\epsilon)>0$ depends only on $\epsilon$.
Proof.
The proof is essentially the same as the proof of [WX22, Proposition 12]. Let $\alpha:[0,\ell]\to X_{g}$ be parameterized by the arc-length parameter such that $\alpha(0)=p$ and $\alpha(\ell)=q$, where $\ell=\ell(\alpha)$. For any $x\in\alpha$, by Lemma 5.1 we have $\operatorname{inj}(x)>\epsilon$. Thus, by standard Sobolev embedding theorem [Tay23],
$\norm{\nabla f}_{L^{\infty}(B(x,\epsilon/2))}\leq\tilde{c}(\epsilon)\sum_{j=0}^{N}\norm{\Delta^{j}(\mathop{}\!\mathrm{d}f)}_{L^{2}(B(x,\epsilon))}\leq\frac{4}{3}\tilde{c}(\epsilon)\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(x,\epsilon))},$
where $\tilde{c}(\epsilon)>0$ only depends on $\epsilon$.
If $\ell<\epsilon/2$, then since $\ell<1$,
(38) |$\absolutevalue{f(p)-f(q)}^{2}$ $\leq\norm{\nabla f}_{L^{\infty}(B(p,\epsilon/2))}^{2}\cdot\ell^{2}\leq\frac{16}{9}\tilde{c}(\epsilon)^{2}\cdot\lambda\cdot\ell.$
If $\ell(\alpha)\geq\epsilon/2$, then for each integer $n\in\{0,\cdots,\lfloor 2\ell/\epsilon\rfloor\}$, set $\alpha_{n}=\alpha(n\epsilon/2)$. We have
(39) |$\absolutevalue{f(p)-f(q)}$ | $\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor-1}\absolutevalue{f(\alpha_{n})-f(\alpha_{n+1})}+\absolutevalue{f(\alpha_{\lfloor 2\ell/\epsilon\rfloor})-f(q)}$ |
$\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\nabla f}_{L^{\infty}(B(\alpha_{n},\epsilon/2))}\cdot\frac{\epsilon}{2}$
$\leq\frac{2}{3}\tilde{c}(\epsilon)\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(B(\alpha_{n},\epsilon))}\cdot\epsilon$
$\leq\frac{2}{3}\tilde{c}(\epsilon)\bigg{(}\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon))}\bigg{)}^{1/2}\bigg{(}\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon^{2}\bigg{)}^{1/2}.$
By Lemma 5.1, $B(\alpha_{n},\epsilon)\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}$, therefore, if $B(\alpha_{n_{1}},\epsilon)\cap B(\alpha_{n_{2}},\epsilon)\neq\emptyset$, then |$\absolutevalue{n_{1}-n_{2}}<2$. Otherwise, we can find a shorter geodesic segment joining $p$ and $q$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, which leads to a contradiction. It then follows that
(40) $\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\norm{\mathop{}\!\mathrm{d}f}^{2}_{L^{2}(B(\alpha_{n},\epsilon))}\leq 2\norm{\mathop{}\!\mathrm{d}f}_{L^{2}(X_{g})}^{2}=2\lambda.$
Moreover, since $\epsilon<1$,
(41) $\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon^{2}\leq\sum_{n=0}^{\lfloor 2\ell/\epsilon\rfloor}\epsilon=(\lfloor 2\ell/\epsilon\rfloor+1)\epsilon\leq 4\ell,$
we deduce from (39), (40), and (41) that
(42) |$\absolutevalue{f(p)-f(q)}^{2}\leq\frac{32}{9}\tilde{c}(\epsilon)^{2}\cdot\lambda\cdot\ell.$
The lemma then follows from (38) and (42) by setting $c(\epsilon)=\frac{32}{9}\tilde{c}(\epsilon)^{2}$. ∎
Similarly to Lemma 3.3, we have
Lemma 5.3.
Let $X_{g}$ be a closed hyperbolic surface, and let $f$ be an eigenfunction of $\Delta$ with eigenvalue $\lambda\leq 1/4-\delta^{2}$. Then
$\int_{X_{g}}f^{2}\mathop{}\!\mathrm{d}A\leq 3\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A,$
where $X^{\mathrm{thick}}_{g}$ is the thick part of $X_{g}$ defined by (36).
Proof.
The proof is essentially the same as that in Lemma 3.3. Since
$\frac{1}{4}-\lambda\geq\delta^{2},$
for any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$, the proof of (8) gives that
$\frac{\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1/\delta)}f^{2}\mathop{}\!\mathrm{d}A}{\int_{\operatorname{\mathcal{C}}(\gamma,w(\gamma))}f^{2}\mathop{}\!\mathrm{d}A}\leq\frac{\int_{0}^{w(\gamma)-1/\delta}\cosh^{2}\!(\delta\rho)\mathop{}\!\mathrm{d}\rho}{\int_{0}^{w(\gamma)}\cosh^{2}\!(\delta\rho)\mathop{}\!\mathrm{d}\rho}\leq\frac{4}{e^{2}},$
thus
$\int_{X_{g}}f^{2}\mathop{}\!\mathrm{d}A$ $=\int_{\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1/\delta)}f^{2}\mathop{}\!\mathrm{d}A+\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A$
$\leq 2\int_{\bigcup_{\gamma\in\mathcal{N}_{\epsilon}(X_{g})}\operatorname{\mathcal{C}}(\gamma,w(\gamma))\setminus\operatorname{\mathcal{C}}(\gamma,w(\gamma)-1/\delta)}f^{2}\mathop{}\!\mathrm{d}A+\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A$
$\leq 3\int_{X^{\mathrm{thick}}_{g}}f^{2}\mathop{}\!\mathrm{d}A.$
The proof is complete. ∎
Now let $\lambda\leq 1/4-\delta^{2}$ be an eigenvalue of $\Delta$ with multiplicity $m(\lambda)$, and let $\{\phi_{i}\}_{1\leq i\leq m(\lambda)}$ be $m(\lambda)$ orthonormal eigenfunctions with eigenvalue $\lambda$. For any $x\in X_{g}$, we define a function $f_{x}$ by
$f_{x}(z):=\frac{\sum_{1\leq i\leq m(\lambda)}\phi_{i}(x)\phi_{i}(z)}{\sqrt{\sum_{1\leq i\leq m(\lambda)}\phi_{i}(x)^{2}}}.$
Then clearly $f_{x}(z)$ is a normalized eigenfunction of $\Delta$ with eigenvalue $\lambda$ satisfying
$f_{x}(x)^{2}=\sum_{i=1}^{m(\lambda)}\phi_{i}(x)^{2}.$
We consider the thick-thin decomposition of $X_{g}$ defined by (36). Inspired by the recent work [GLN24] of Gross-Lachman-Nachmias, for any $j\geq 0$, we also set
$\mu_{j}=2^{j}\sqrt{\lambda}\text{\quad and\quad}R_{j}=c_{1}\cdot 2^{j}/\sqrt{\lambda}.$
The constant $c_{1}>0$, depending only on $\epsilon$, will be determined in Lemma 5.6. Now fix some $k\in[1,I_{\epsilon}]$. We first give two definitions.
Definition 5.4.
For each $j\geq 0$, define
$G_{j}:=\Big{\{}x\in\mathcal{A}_{k}:\ \mu_{j-1}\leq f_{x}(x)^{2}<\mu_{j}\Big{\}},$
where we set $\mu_{-1}=0$.
Definition 5.5.
For any $x\in\mathcal{A}_{k}$ and any $R>0$, define
$B_{x,R}:=\left\{z\in\mathcal{A}_{k}\cup\mathcal{S}_{k}:\ \begin{gathered}\text{there exists a geodesic segment of length $\leq R$ joining}\\ \text{$x$ and $z$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$}\end{gathered}\right\}.$
We have the following lemmas.
Lemma 5.6.
Suppose $x\in G_{j}$. If $j\geq 1$, then
$\operatorname{Area}(B_{x,R_{j}})\leq\frac{4}{\mu_{j-1}}.$
Proof.
We first prove that
(43) $B_{x,R_{j}}\subset\Big{\{}z\in X_{g}:\ f_{x}(z)\geq\frac{f_{x}(x)}{2}\Big{\}}.$
Otherwise, suppose that there would exist some $z_{0}\in B_{x,R_{j}}$ such that
$f_{x}(z_{0})<\frac{f_{x}(x)}{2}.$
By the definition of $B_{x,R_{j}}$, we can find a shortest geodesic segment of length $\leq R_{j}$ joining $x$ and $z_{0}$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. Thus, by Lemma 5.2 we have
$\frac{\mu_{j-1}}{4}\leq\frac{f_{x}(x)^{2}}{4}\leq\absolutevalue{f_{x}(x)-f_{x}(z_{0})}^{2}\leq c_{2}\cdot R_{j}\cdot\lambda,$
where $c_{2}=c(\epsilon)$ depends only on $\epsilon$. It follows that
$1\leq 8c_{2}\cdot c_{1}.$
If we set $c_{1}=1/(9c_{2})$, then it leads to a contradiction. Thus, it follows from (43) and the Markov inequality that for any $x\in G_{j}$,
$\operatorname{Area}(B_{x,R_{j}})\leq\operatorname{Area}\left(\Big{\{}z\in X_{g}:\ f_{x}(z)\geq\frac{f_{x}(x)}{2}\Big{\}}\right)\leq\frac{\int_{X_{g}}f_{x}^{2}(z)}{\left(\frac{f_{x}(x)}{2}\right)^{2}}\leq\frac{4}{\mu_{j-1}}.$
This completes the proof. ∎
We will use a different notion of diameter which is defined as follows.
Definition 5.7.
$\operatorname{Diam}_{c}(\mathcal{A}_{k}):=\sup_{x,z\in\mathcal{A}_{k}}\left\{\operatorname{\ell}(\alpha):\ \begin{gathered}\text{$\alpha$ is a shortest geodesic segment joining $x$ and $z$}\\ \text{that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$}\end{gathered}\right\}.$
Lemma 5.8.
Suppose $x\in\mathcal{A}_{k}$ and $R>2c_{1}$. If $R<\operatorname{Diam}_{c}(\mathcal{A}_{k})$, then
$\operatorname{Area}(B_{x,R/2})\geq c_{3}\cdot R,$
where $c_{3}>0$ depends only on $\epsilon$.
Proof.
Since $R<\operatorname{Diam}_{c}(\mathcal{A}_{k})$, by definition there exist some $z\in\mathcal{A}_{k}$ and a shortest geodesic segment $\alpha$ from $x$ to $z$ of length $R/2$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$. By Lemma 5.1, we know that $\alpha\subset\mathcal{A}_{k}\cup\mathcal{S}_{k}$, hence
$\alpha\subset B_{x,R/2}.$
Let $\alpha:[0,R/2]\to X_{g}$ be a geodesic of arc-length parameter with $\alpha(0)=x$. Then for each integer $n\in\{0,\cdots,\lfloor R/c_{1}\rfloor-1\}$, we have
$B(\alpha(nc_{1}/2),c_{1}/4)\subset B_{x,R/2},$
and the geodesic balls $B(\alpha(nc_{1}/2),c_{1}/4)$, $11\leq n\leq\lfloor R/c_{1}\rfloor-$ 1, are pairwise disjoint. Moreover, by Lemma 5.1 we have that for all $10\leq n\leq\lfloor R/c_{1}\rfloor-$ 1,
$\operatorname{inj}(\alpha(nc_{1}/2))>\epsilon.$
Therefore,
$\operatorname{Area}(B_{x,R/2})$ $\geq\sum_{n=0}^{\lfloor R/c_{1}\rfloor-1}\operatorname{Area}\big{(}B(\alpha(nc_{1}/2),c_{1}/4)\big{)}$
$\geq\min\Big{\{}\pi\epsilon^{2},\frac{\pi c_{1}^{2}}{16}\Big{\}}\cdot\lfloor R/c_{1}\rfloor$
$\geq c_{3}\cdot R,$
where $c_{3}$ depends only on $\epsilon$. The proof is complete. ∎
Now we give an upper bound of the integral of $f_{x}(x)^{2}$ over $\mathcal{A}_{k}$.
Lemma 5.9.
There exists some constant $C>0$ depending only on $\epsilon$ such that
[…]
Proof.
Let $j_{0}\geq 0$ be the maximal index such that $G_{j_{0}}\neq\emptyset$ (see Definition 5.4). If $j_{0}=0$, then
(44) $\int_{\mathcal{A}_{k}}f_{x}(x)^{2}\leq\mu_{0}\cdot\operatorname{Area}(G_{0})=\sqrt{\lambda}\cdot\operatorname{Area}(\mathcal{A}_{k}).$
From now on, we always assume that $j_{0}\geq 1$. Note that
(45) $\int_{\mathcal{A}_{k}}f_{x}(x)^{2}\leq\sum_{j=0}^{j_{0}}\mu_{j}\cdot\operatorname{Area}(G_{j}).$
Case-1. $R_{j_{0}}\geq\operatorname{Diam}_{c}(\mathcal{A}_{k})$. Then by definition we know that for any $x\in G_{j_{0}}$
$\mathcal{A}_{k}\subset B_{x,R_{j_{0}}}.$
Since $j_{0}\geq 1$, it follows from Lemma 5.6 that
$\operatorname{Area}(\mathcal{A}_{k})\leq\operatorname{Area}(B_{x,R_{j_{0}}})\leq\frac{4}{\mu_{j_{0}-1}},$
we thus have
$\mu_{0}<\mu_{1}<\cdots<\mu_{j_{0}}\leq\frac{8}{\operatorname{Area}(\mathcal{A}_{k})}.$
By the definition of $G_{j}$ we thus have
(46) $\sum_{j=0}^{j_{0}}\mu_{j}\cdot\operatorname{Area}(G_{j})\leq\frac{8}{\operatorname{Area}(\mathcal{A}_{k})}\sum_{j=0}^{j_{0}}\operatorname{Area}(G_{j})=8.$
Case-2. $R_{j_{0}}<\operatorname{Diam}_{c}(\mathcal{A}_{k})$. For any $1\leq j\leq j_{0}$, let $\{B_{x_{i},R_{j}/2}\}_{1\leq i\leq l}$ be a maximal collection of disjoint sets such that $x_{i}\in G_{j}$, $\operatorname{\forall}1\leq i\leq l$. By Lemma 5.8 we have
(47) $l\leq\frac{\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k})}{c_{3}R_{j}}.$
Next we claim that
(48) $G_{j}\subset\bigcup_{1\leq i\leq l}B_{x_{i},R_{j}}.$
Otherwise, there would exist some $x\in G_{j}$ such that any geodesic segment joining $x$ and $x_{i}$ that does not intersect with any $\gamma\in\mathcal{N}_{\epsilon}(X_{g})$ has length $\geq R_{j}$. Thus, $B_{x,R_{j}/2}$ is disjoint with any $B_{x_{i},R_{j}/2}$, leading to a contradiction because our choice of $\{B_{x_{i},R_{j}/2}\}_{1\leq i\leq l}$ is maximal. Combining (48), Lemma 5.6, Lemma 5.8 and (47), we deduce that
$\operatorname{Area}(G_{j})\leq\sum_{i=1}^{l}\operatorname{Area}(B_{x_{i},R_{j}})\leq\sum_{i=1}^{l}\frac{4}{\mu_{j-1}}\leq\frac{\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k})}{c_{3}R_{j}}\frac{4}{\mu_{j-1}}.$
Therefore, we have
(49) $\sum_{j=0}^{j_{0}}\mu_{j}\cdot\operatorname{Area}(G_{j})$ $\leq\mu_{0}\cdot\operatorname{Area}(G_{0})+\sum_{j=1}^{j_{0}}\mu_{j}\cdot\frac{4\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k})}{c_{3}\mu_{j-1}R_{j}}$
$\leq\sqrt{\lambda}\cdot\operatorname{Area}(\mathcal{A}_{k})+\sum_{j=1}^{j_{0}}\frac{8\sqrt{\lambda}\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k})}{c_{1}c_{3}2^{j}}$
$\leq C\sqrt{\lambda}\cdot\operatorname{Area}(\mathcal{A}_{k}\cup\mathcal{S}_{k}).$
The lemma then follows from (44), (45), (46) and (49). ∎
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
By Lemma 5.3 and Lemma 5.9 we have
$m(\lambda)=\int_{X_{g}}\sum_{i=1}^{m(\lambda)}\phi_{i}(x)^{2}$ $\leq 3\int_{X^{\mathrm{thick}}_{g}}\sum_{i=1}^{m(\lambda)}\phi_{i}(x)^{2}$
$=3\sum_{k=1}^{I_{\epsilon}}\int_{\mathcal{A}_{k}}f_{x}(x)^{2}$
[…]
$\leq 3C\sqrt{\lambda}\cdot\operatorname{Area}(X_{g})+24I_{\epsilon}.$
The proof is then finished by Gauss-Bonnet. ∎